# Retrieval-Augmented Generation with Conflicting Evidence

### v1

@@ Acknowledgments
This work was supported by NSF-CAREER Award 1846185, DARPA ECOLE Program No. HR00112390060, the Microsoft Accelerate Foundation Models Research (AFMR) grant program, and NSF-AI Engage Institute DRL-2112635. Any opinions, findings, and conclusions or recommendations in this work are those of the author(s) and do not necessarily reflect the views of the sponsors.
@@ Appendix A RamDocs: Dataset Statistics
Appendix B Astute RAG
@@
Appendix C Prompts
@@
Multi-Agent (Individual Agent) You are an agent reading a document to answer a question. Question: {question} Document: {document} The following responses are from other agents as additional information. # Optional {history} # Optional Answer the question based on the document and other agents’ responses. Provide your answer and a step-by-step reasoning explanation. Please follow the format: ’Answer: {}. Explanation: {}.
@@
Multi-Agent (Aggregator) You are an aggregator reading answers from multiple agents. If there are multiple answers, please provide all possible correct answers and also provide a step-by-step reasoning explanation. If there is no correct answer, please reply ’unknown’. Please follow the format: “All Correct Answers: []. Explanation: .” The following are examples: Question: In which year was Michael Jordan born? Agent responses: Agent 1: Answer: 1963. Explanation: The document clearly states that Michael Jeffrey Jordan was born on February 17, 1963. Agent 2: Answer: 1956. Explanation: The document states that Michael Irwin Jordan was born on February 25, 1956. However, it’s important to note that this document seems to be about a different Michael Jordan, who is an American scientist, not a basketball player. The other agents’ responses do not align with the information provided in the document. Agent 3: Answer: 1998. Explanation: According to the document provided, Michael Jeffrey Jordan was born on February 17, 1998. Agent 4: Answer: Unknown. Explanation: The provided document focuses on Jordan’s college and early professional career, mentioning his college championship in 1982 and his entry into the NBA in 1984, but it does not include information about his birth year. All Correct Answers: [”1963”, ”1956”]. Explanation: Agent 1 is talking about the basketball player Michael Jeffrey Jordan, who was born on February 17, 1963, so 1963 is correct. Agent 2 is talking about another person named Michael Jordan, who is an American scientist, and he was born in 1956. Therefore, the answer 1956 from Agent 2 is also correct. Agent 3 provides an error stating Michael Jordan’s birth year as 1998, which is incorrect. Based on the correct information from Agent 1, Michael Jeffrey Jordan was born on February 17, 1963. Agent 4 does not provide any useful information. Question: {question} Agent responses: {agent responses list}
@@
Astute RAG (Adaptive Passage Generation) Generate a document that provides accurate and relevant information to answer the given question. If the information is unclear or uncertain, explicitly state ’I don’t know’ to avoid any hallucinations. Question: {question} Document:
@@
Astute RAG (Iterative Knowledge Consolidation) Task: Consolidate information from both your own memorized documents and externally retrieved documents in response to the given question. * For documents that provide consistent information, cluster them together and summarize the key details into a single, concise document. * For documents with conflicting information, separate them into distinct documents, ensuring each captures the unique perspective or data. * Exclude any information irrelevant to the query. For each new document created, clearly indicate: * Whether the source was from memory or an external retrieval. * The original document numbers for transparency. Initial Context: {context_init} Last Context: {context} Question: {question} New Context:
@@
Astute RAG (Knowledge Consolidation and Answer Finalization) Task: Answer a given question using the consolidated information from both your own memorized documents and externally retrieved documents. Step 1: Consolidate information * For documents that provide consistent information, cluster them together and summarize the key details into a single, concise document. * For documents with conflicting information, separate them into distinct documents, ensuring each captures the unique perspective or data. * Exclude any information irrelevant to the query. For each new document created, clearly indicate: * Whether the source was from memory or an external retrieval. * The original document numbers for transparency. Step 2: Propose Answers and Assign Confidence For each group of documents, propose a possible answer and assign a confidence score based on the credibility and agreement of the information. Step 3: Select all Correct Answers After evaluating all groups, select all accurate and well-supported answers. Initial Context: {context_init} Consolidated Context: {context} # optional Question: {question} Answer:

### v2

@@ Acknowledgments
We would like to thank the anonymous reviewers for their feedback. This work was supported by NSF-CAREER Award 1846185, DARPA ECOLE Program No. HR00112390060, the Microsoft Accelerate Foundation Models Research (AFMR) grant program, and NSF-AI Engage Institute DRL-2112635. Any opinions, findings, and conclusions or recommendations in this work are those of the author(s) and do not necessarily reflect the views of the sponsors.
@@ Appendix A RamDocs: Dataset Statistics
Appendix B Speculative RAG
Speculative RAG (Wang et al., 2025) first clusters the retrieved documents into $k$ clusters based on their perspectives using an embedding model, InBedder-RoBERTa-large (Peng et al., 2024). From each cluster, it samples one document to form a subset $\delta_{j}$, repeating this process until unique subsets are created. Each subset $\delta_{j}$ is fed into a specialist RAG drafter [^3] As (at time of publication) Wang et al. (2025)’s drafter model is not publicly available, we use different models listed in Table 1 instead. to generate a draft answer $\alpha_{j}$ and its rationale $\beta_{j}$. Since our setting may have multiple correct answers, we pass all answers and their rationales to the aggregator to generate the final answer.
Appendix C Astute RAG
@@
Appendix D Computational Efficiency of Madam-RAG
To improve efficiency, we have incorporated an early stopping criterion in Madam-RAG: if all agents retain their answers across consecutive rounds, the debate terminates early. In practice, we observe convergence within 1-2 rounds for most examples. To provide a concrete comparison, we calculated the average number of input and output tokens for three representative methods: prompt-based, single agent with self-reflection, and MADAM-RAG.
Method Avg Input Tokens Avg Output Tokens
Prompt-based 822.69 126.25
Single Agent with Self-reflection 5033.92 809.10
MADAM-RAG (Ours) 3186.45 1547.18
Table 2: Average input and output token usage of different methods.
As shown in Table 2, MADAM-RAG requires 3186 input tokens and 1547 output tokens on average—higher than prompt-based RAG (822/126) but comparable to the Single Agent with Self-reflection baseline when taking both input and output tokens into account. This demonstrates that while MADAM-RAG introduces some overhead due to its multi-agent setup, it remains computationally tractable and roughly as efficient as strong single-agent iterative alternatives, while outperforming both zero-shot and single-agent approaches.
Appendix E Prompts
@@
Single Agent with Self-reflection Initial Generation:
You are an expert in retrieval question answering.
You will be provided a question with multiple documents. Please answer the question based on the documents. If there are multiple answers, please provide all possible correct answers and also provide a step-by-step reasoning explanation. If there is no correct answer, please reply ’unknown’. Please follow the format: ’All Correct Answers: []. Explanation: {}.’ The following are examples: Question: In which year was Michael Jordan born? Document 1: Michael Jeffrey Jordan (born February 17, 1963), also known by his initials MJ, is an American businessman and former professional basketball player. He played 15 seasons in the National Basketball Association (NBA) between 1984 and 2003, winning six NBA championships with the Chicago Bulls. He was integral in popularizing basketball and the NBA around the world in the 1980s and 1990s, becoming a global cultural icon. Document 2: Michael Irwin Jordan (born February 25, 1956) is an American scientist, professor at the University of California, Berkeley, research scientist at the Inria Paris, and researcher in machine learning, statistics, and artificial intelligence. Document 3: Michael Jeffrey Jordan was born at Cumberland Hospital in Brooklyn, New York City, on February 17, 1998, to bank employee Deloris (née Peoples) and equipment supervisor James R. Jordan Sr. He has two older brothers, James Jr. and Larry, as well as an older sister named Deloris and a younger sister named Roslyn. Jordan and his siblings were raised Methodist. Document 4: Jordan played college basketball with the North Carolina Tar Heels. As a freshman, he was a member of the Tar Heels’ national championship team in 1982. Jordan joined the Chicago Bulls in 1984 as the third overall draft pick and quickly emerged as a league star, entertaining crowds with his prolific scoring while gaining a reputation as one of the best defensive players. All Correct Answers: [”1963”, ”1956”]. Explanation: Document 1 is talking about the basketball player Michael Jeffrey Jordan, who was born on Februray 17, 1963, so 1963 is correct. Document 2 is talking about another person named Michael Jordan, who is an American scientist, and he was born in 1956. Therefore, the answer 1956 from Document 2 is also correct. Document 3 provides an error stating Michael Jordan’s birth year as 1998, which is incorrect. Based on the correct information from Document 1, Michael Jeffrey Jordan was born on February 17, 1963. Document 4 does not provide any useful information. Question: {question} {context}
Review Generation:
You are an expert in retrieval question answering.
You will be provided a question with multiple documents. Please answer the question based on the documents. If there are multiple answers, please provide all possible correct answers and also provide a step-by-step reasoning explanation. If there is no correct answer, please reply ’unknown’. Question: {question} {context} {answer} Review your previous answer and find problems with your answer.
Refine Generation:
You are an expert in retrieval question answering.
You will be provided a question with multiple documents. Please answer the question based on the documents. If there are multiple answers, please provide all possible correct answers and also provide a step-by-step reasoning explanation. If there is no correct answer, please reply ’unknown’. Question: {question} {context} {answer} Review your previous answer and find problems with your answer. {review} Based on the problems you found, improve your answer. Please reiterate your answer with the format: ’All Correct Answers: []. Explanation: {}.’
Speculative RAG Drafter:
Answer the question based on the documents. Also provide rationale for your response.
Question: {question} {documents} Please follow the format: Response: {}. Rationale: {}. You are an expert in question answering.
Aggregator:
You are an aggregator reading answers from multiple responses.
@@
If there are multiple answers, please provide all possible correct answers and also provide a step-by-step reasoning explanation. If there is no correct answer, please reply ’unknown’. Please follow the format: ’All Correct Answers: []. Explanation: {}.’ The following are examples: Question: In which year was Michael Jordan born? Responses: Response 1: Response: 1963. Rationale: The document clearly states that Michael Jeffrey Jordan was born on February 17, 1963. Response 2: Response: 1956. Rationale: The document states that Michael Irwin Jordan was born on February 25, 1956. However, it’s important to note that this document seems to be about a different Michael Jordan, who is an American scientist, not the basketball player. The other agents’ responses do not align with the information provided in the document. Response 3: Response: 1998. Rationale: The According to the document provided, Michael Jeffrey Jordan was born on February 17, 1998. Response 4: Response: Unknown. Rationale: The provided document focuses on Jordan’s college and early professional career, mentioning his college championship in 1982 and his entry into the NBA in 1984, but it does not include information about his birth year. All Correct Answers: [”1963”, ”1956”]. Explanation: Response 1 is talking about the basketball player Michael Jeffrey Jordan, who was born on Februray 17, 1963, so 1963 is correct. Response 2 is talking about another person named Michael Jordan, who is an American scientist, and he was born in 1956. Therefore, the answer 1956 from Response 2 is also correct. Response 3 provides an error stating Michael Jordan’s birth year as 1998, which is incorrect. Based on the correct information from Response 1, Michael Jeffrey Jordan was born on February 17, 1963. Response 4 does not provide any useful information. Question: {query} Responses: {all drafted responses}
Astute RAG Adaptive Passage Generation:
@@
Generate a document that provides accurate and relevant information to answer the given question. If the information is unclear or uncertain, explicitly state ’I don’t know’ to avoid any hallucinations.
Question: {question} Document:
Iterative Knowledge Consolidation:
Task: Consolidate information from both your own memorized documents and externally retrieved documents in response to the given question.
@@
* For documents that provide consistent information, cluster them together and summarize the key details into a single, concise document. * For documents with conflicting information, separate them into distinct documents, ensuring each captures the unique perspective or data. * Exclude any information irrelevant to the query. For each new document created, clearly indicate: * Whether the source was from memory or an external retrieval. * The original document numbers for transparency. Initial Context: {context_init} Last Context: {context} Question: {question} New Context:
Knowledge Consolidation and Answer Finalization:
Task: Answer a given question using the consolidated information from both your own memorized documents and externally retrieved documents.
@@
Step 1: Consolidate information * For documents that provide consistent information, cluster them together and summarize the key details into a single, concise document. * For documents with conflicting information, separate them into distinct documents, ensuring each captures the unique perspective or data. * Exclude any information irrelevant to the query. For each new document created, clearly indicate: * Whether the source was from memory or an external retrieval. * The original document numbers for transparency. Step 2: Propose Answers and Assign Confidence For each group of documents, propose a possible answer and assign a confidence score based on the credibility and agreement of the information. Step 3: Select all Correct Answers After evaluating all groups, select all accurate and well-supported answers. Initial Context: {context_init} Consolidated Context: {context} # optional Question: {question} Answer:
MADAM-RAG Individual Agent:
You are an agent reading a document to answer a question.
Question: {question} Document: {document} The following responses are from other agents as additional information. # Optional {history} # Optional Answer the question based on the document and other agents’ responses. Provide your answer and a step-by-step reasoning explanation. Please follow the format: ’Answer: {}. Explanation: {}.
Aggregator:
You are an aggregator reading answers from multiple agents.
If there are multiple answers, please provide all possible correct answers and also provide a step-by-step reasoning explanation. If there is no correct answer, please reply ’unknown’. Please follow the format: “All Correct Answers: []. Explanation: .” The following are examples: Question: In which year was Michael Jordan born? Agent responses: Agent 1: Answer: 1963. Explanation: The document clearly states that Michael Jeffrey Jordan was born on February 17, 1963. Agent 2: Answer: 1956. Explanation: The document states that Michael Irwin Jordan was born on February 25, 1956. However, it’s important to note that this document seems to be about a different Michael Jordan, who is an American scientist, not a basketball player. The other agents’ responses do not align with the information provided in the document. Agent 3: Answer: 1998. Explanation: According to the document provided, Michael Jeffrey Jordan was born on February 17, 1998. Agent 4: Answer: Unknown. Explanation: The provided document focuses on Jordan’s college and early professional career, mentioning his college championship in 1982 and his entry into the NBA in 1984, but it does not include information about his birth year. All Correct Answers: [”1963”, ”1956”]. Explanation: Agent 1 is talking about the basketball player Michael Jeffrey Jordan, who was born on February 17, 1963, so 1963 is correct. Agent 2 is talking about another person named Michael Jordan, who is an American scientist, and he was born in 1956. Therefore, the answer 1956 from Agent 2 is also correct. Agent 3 provides an error stating Michael Jordan’s birth year as 1998, which is incorrect. Based on the correct information from Agent 1, Michael Jeffrey Jordan was born on February 17, 1963. Agent 4 does not provide any useful information. Question: {question} Agent responses: {agent responses list}